OpenStudy (anonymous):

how to find the area of an equilateral triangle with a side length of 10.4 meters

OpenStudy (anonymous):

http://www.ezscrap.net/EquilateralTriangle_1000.gif Bear this triangle in mind^. The area of a triangle = 0.5bh b=a=10.7 m h=sqrt(a²-(0.5a)²)=sqrt(0.75a²)=sqrt(0.75*10.4²)=9.007 So: Area=0.5bh=0.5*10.4*9.007=46.837 m²

OpenStudy (jit4won):

area of an equilateral triangle = √ 3 / 4 ( a )^2 where a = side of the equilateral triangle now you can say area = √ 3 / 4 ( 10.4) ^ 2 = √ 3 / 4 x 10.4 x 10.4 =27.04√ 3

OpenStudy (anonymous):

first find the heigh 10.4/2=5.2 \[\sqrt{10.4^{2}-5.2^{2}} =\sqrt{108.16-27.4}\] the heigh is:\[\sqrt{80.76}\] A=(b*h)/2 \[(10.4*\sqrt{80.76})\div 2\] A= 46.730